OpenStudy (anonymous):

The rabbit population on a small island is given by the function below, where t is the time in months since observation of the rabbit population began. P(t)=110t-0.4t^4+1000 What is the maximum population? Please show work.

hero (hero):

Well, you could simply graph the function to find it's max value.

hero (hero):

You could also take the derivative of the function too.

OpenStudy (a_clan):

find the maxima

hero (hero):

John, do you still need help with this?

OpenStudy (anonymous):

equation rewrite p(t)= -.4t^4 + 110t +1000 p'(t)= -1.6t^3 + 110 maximum and min is where this equation = 0 0=-1.6t^3 + 110 1.6t^3=110 t^3 =68.75 t=4.0966

OpenStudy (anonymous):

looks about right in graph form

hero (hero):

hobeau, you found the point where the slope of p(t) = 0. You didn't find the actual max.